4. If fixed costs are $100 and variable costs are $200 at an output level of 30 units, what are the average fixed costs, average variable costs, and average total costs?

*Please show all work this time.

Quantity

Fixed Cost

Total Cost

Average Fixed Cost

Average Variable Cost

Average Total Cost

Marginal Cost

10

5,000

5,500

20

5,000

5,800

30

5,000

6,000

40

5,000

6,100

50

5,000

6,150

60

5,000

6,175
Answer #1

Ans.4. fixed costs (FC) $100 and variable costs (VC) $200 at an output level (Q) of 30 units,

Total cost (TC) = FC + VC

TC = $100 + $200

TC = $300

i) Average Fixed Cost (AFC ) = FC /Q

AFC = $100/30

AFC = $3.33

Average Fixed Cost is $3.33.

ii) Average Variable Cost (AVC ) = VC / Q

AVC = $200/30

AVC = $6.67

Average Variable Cost is $6.67.

iii) Average Total Cost (ATC ) = TC/ Q

ATC = $300/30

ATC = $10

Average Total Cost is $10.

Quantity Fixed Cost Total Cost Variable Cost Average Fixed Cost Average Variable Cost Average Total Cost Marginal Cost
10 5,000 5,500 500 500 50 550 50
20 5,000 5,800 800 250 40 290 30
30 5,000 6,000 1,000 166.67 33.33 200 20
40 5,000 6,100 1,100 125 27.50 152.50 10
50 5,000 6,150 1,150 100 23 123 5
60 5,000 6,175 1,175 83.33 19.58 102.92 2.5

Working Note:

1) TC = FC + VC

2) VC = TC - FC

3) AFC = FC/Q

4) AVC = VC/Q

5) ATC = TC/Q

6) MC = change in TC/ change in Q

7) At Q = 0, TC = $5,000